# ZeroDVFS: Zero-Shot LLM-Guided Core and Frequency Allocation for Embedded Platforms

### v1

@@ 4.2 Environment Design and Modeling
Different Environment Models and their Computational Complexity. To determine the most efficient architecture for the environment model, this study evaluates several state-of-the-art regression models in predicting profiler states ($s^{\prime}_{\text{profiler}}$) and temperature states ($s^{\prime}_{\text{temp}}$). The models are trained using data that includes profiler state information ($s_{\text{profiler}}$), temperature state ($s_{\text{temp}}$), and their corresponding actions ($a_{\text{profiler}}$ for the profiler model and $a_{\text{temp}}$ for the temperature model). The regression models considered are Simple Fully Connected Networks (FCNs), Convolutional Neural Networks (CNNs), Recurrent Neural Networks (RNNs), Long Short-Term Memory Networks (LSTMs), and Attention-Based Networks. Model selection prioritizes both prediction accuracy and inference latency to enable efficient planning.
$\text{Parameters}_{\text{FCN}}$ $=N_{\text{input}}\times N_{\text{hidden}}$
$\quad+N_{\text{hidden}}\times N_{\text{state}}^{\prime}$
$\quad+N_{\text{hidden}}+N_{\text{state}}^{\prime}$
FCN models achieve inference latencies of 2-5ms on embedded platforms.
$\quad+C_{\text{out}}$
CNN inference latency ranges from 4-8ms.
$\text{Parameters}_{\text{RNN}}$ $=N_{\text{input}}\times N_{\text{hidden}}$
$\quad+N_{\text{hidden}}\times N_{\text{hidden}}$
$\quad+N_{\text{hidden}}+N_{\text{hidden}}\times N_{\text{state}}^{\prime}$
$\quad+N_{\text{state}}^{\prime}$
Long Short-Term Memory Networks extend RNNs by incorporating gates to better manage long-term dependencies. An LSTM with $N_{\text{hidden}}$ neurons has approximately three times the number of parameters compared to a standard RNN due to the additional forget, input, and output gates, resulting in higher inference latency.
Attention mechanisms exhibit higher inference latency but provide superior prediction accuracy.
@@ 4.3 Cross-Platform Model Transfer and Adaptation
4.3.1 Transfer Learning Methodology
We employ a two-stage transfer learning approach:
@@ 4.3.1 Transfer Learning Methodology
4.3.2 Feature Classification for Transfer
We categorize features based on their transferability:
@@ 4.3.2 Feature Classification for Transfer
4.3.3 Transfer Learning Results
The cross-platform transfer learning results are presented in Section 5 (Table 5). Our experiments evaluate zero-shot transfer from Jetson TX2 to Orin NX and RubikPi platforms, demonstrating the domain shift effects when deploying learned models across different embedded architectures. The source platform achieves 10.9% MAPE (Mean Absolute Percentage Error), while zero-shot transfer shows 64.5% and 73.2% MAPE for Orin NX and RubikPi respectively, reflecting significant architectural differences between platforms.
4.3.4 Comparison with Table-Based Approaches
@@ 4.4 LLM-Based Semantic Feature Extraction
4.4.1 Motivation: From Syntax to Semantics
The challenge of workload characterization without execution traces requires distinguishing between what code contains versus how it behaves. Traditional static analysis tools such as Tree-sitter [57] and compiler front-ends excel at extracting syntactic properties, such as counting loops, identifying OpenMP pragmas, and measuring code complexity metrics. However, these tools fundamentally cannot infer semantic properties that determine runtime behavior.
@@ 4.4.1 Motivation: From Syntax to Semantics
4.4.2 Two-Stage Feature Extraction Pipeline
Our feature extraction operates in two complementary stages. The first stage employs Tree-sitter to extract syntactic features capturing the structural composition of OpenMP programs: control flow metrics such as loop depth and nesting complexity, OpenMP directive counts including parallel regions and task constructs, synchronization primitives such as critical sections and atomic operations, and variable scope classifications distinguishing shared, private, and reduction variables. These 17 features provide a syntactic fingerprint of code structure but cannot distinguish semantically different algorithms with similar structural properties.
@@ 4.4.2 Two-Stage Feature Extraction Pipeline
4.4.3 Multi-Model Agreement and Reliability
Extracting semantic features through LLM inference raises natural questions about reliability and consistency. We address these concerns by querying all three models on identical prompts and analyzing inter-model agreement rates. Table 1 presents pairwise and unanimous agreement percentages across the 42 benchmarks in our evaluation suite.
@@ 4.4.3 Multi-Model Agreement and Reliability
4.4.4 Cost-Effective Zero-Shot Prediction
The practical viability of LLM-based feature extraction depends critically on its cost relative to traditional profiling. Our analysis shows favorable cost trade-offs. Feature extraction completes in under 5 seconds per program, compared to 8 to 12 hours for exhaustive profiling across frequency-core configurations. The monetary cost ranges from \$0.0015 per program using DeepSeek-V3 alone to \$0.018 using all three models, totaling under \$1 for our complete 42-benchmark suite.
@@ 5.1 Implemented Algorithms
Model-Free Single-Agent RL (zTT). We evaluate a single-agent, model-free RL approach implemented on the Jetson TX2 platform, as described in [30]. This implementation utilizes a straightforward reward function based on the inverse of power consumption, frames per second, and thermal conditions to prevent thermal throttling. Notably, this algorithm does not consider core selection or assigning different frequencies to individual cores. Although a more recent extension of zTT is introduced in [38], we rely on the original implementation due to the availability of code, which allows us to compare the performance of different models effectively. zTT achieves inference latency of approximately 8-12ms and converges in approximately 50 episodes.
Model-Based Single-Agent RL (DynaQ). We adapted the Dyna-Q algorithm for a multi-core processor environment, similar to Algorithm 1, excluding the thermal agent. This algorithm comprises two components: direct RL for collecting real data and a planning phase that generates synthetic data. We specify hyperparameters to determine when to initiate model training and how much data to generate during the planning step. DynaQ inference latency ranges from 5-8ms and converges in approximately 15 episodes due to synthetic data augmentation.
Generative Model-Based Single-Agent RL (PlanGAN). We adopted the PlanGAN algorithm [11], which integrates Generative Adversarial Networks (GANs) with model-based RL. This approach enhances planning and policy learning by generating realistic synthetic trajectories. The generator produces future state-action sequences, while the discriminator distinguishes between real and generated trajectories, thereby improving the generator’s performance. This allows the agent to simulate future outcomes and plan actions without constant interaction with the environment, reducing data inefficiency and enhancing decision-making. PlanGAN has higher inference latency (12-18ms) due to GAN forward passes.
Multi-Agent Model-Free (MAMF) RL. We introduce a multi-agent model-free RL approach, as outlined in Algorithm 1, to define core selection actions based on thermal and profiling data. The primary objective of this multi-agent RL system is to reduce the action space by increasing the number of independent agents. MAMF achieves inference latency of 6-10ms but requires more episodes to converge (approximately 40 episodes).
Multi-Agent Model-Based (MAMB) RL. This approach follows the same architecture as described in Algorithm 1 and uses a similar agent definition to the model-free method. It consists of two agents: the profiler agent and the thermal agent, and an environment model. The environment model is trained based on predefined hyperparameters and generates synthetic data accordingly. MAMB achieves inference latency of 5-8ms and converges in 25 episodes.
MAMB RL with D3QN Agents. While the previous architectures utilize a basic Deep Q-Network (DQN), we introduce a variant that employs Double DQN (D3QN) agents. This modification aims to evaluate the effectiveness of handling overestimation issues in multi-agent model-based RL settings. MAMBRL D3QN maintains similar inference latency (5-9ms) while achieving fastest convergence at 20 episodes with superior Q-value stability.
@@ 5.2 Experimental Platform, Benchmark, and Evaluation
Experimental Platforms: The algorithms are implemented on an environment inspired from zTT for a better evaluation. The Linux kernel tools are used to profile and evaluate the application that is tested on embedded processor Jetson TX2, Jetson Orin NX, RubikPi, and superscalar 4-core Core I7 8th gen. Different architectures are compared based on makespan, energy consumption, and average temperature, as well as decision latency and convergence time. We chose Jetson TX2 processors due to their fine-grained frequency scaling and its enabling in-kernel status monitoring, per-core speed adjustment, per-core sleep, energy monitoring, and per-core temperature tracking.
Table 2: Experimental Platform Specifications
Specification Jetson TX2 Orin NX RubikPi
Architecture ARM Cortex-A57 ARM Cortex-A78AE Qualcomm Kryo
+ Denver 2 + Cortex-A78AE 585
CPU Cores 6 8 8
Thermal Zones 8 9 36
Table 2 summarizes the key specifications of the embedded platforms used in our evaluation. The Jetson TX2 serves as our primary evaluation target with 6 CPU cores (4 ARM Cortex-A57 + 2 Denver 2) and 8 thermal zones. The Orin NX provides 8 ARM Cortex-A78AE cores with 9 thermal zones, representing next-generation NVIDIA embedded platforms. RubikPi features 8 Qualcomm Kryo 585 cores with 36 thermal zones, enabling evaluation of cross-vendor transfer within the ARM ecosystem.
Targeted Benchmark: All the architectures are trained on the Barcelona OpenMP Tasks (BOTs), and CPU affinity is applied to allocate tasks to the corresponding core [20]. We tested and evaluated multiple benchmarks; this paper primarily reports results for FFT and Strassen due to their task dependencies and parallel execution patterns representative of AI workloads. Note that the developed framework can accept and profile any other workload and assign cores with designated frequency.
Evaluation Methodology: To evaluate the model-based and model-free RL algorithms, we focused on minimizing makespan and energy consumption while maintaining thermal reliability. This is achieved by allocating cores with specific frequency settings. Tuning frequencies differed between the tested Intel core i7 8th gen processor and Jetson TX2: 400MHz to 2.1GHz for the Core i7 8650 and 345MHz to 2.035GHz in Jetson TX2.
The experiments on Jetson TX2 ran on Ubuntu 18.04 and on Intel Core i7 ran on Ubuntu 22.04 with the FIFO real time scheduler with high priority [47] from Preemptible Linux kernel version 4.9.337. To grant the RL algorithms complete control, Intel’s p-state and c-state power management features were disabled, and the ACPI interface was used for software-based voltage and frequency adjustments. Additionally, the open-source ACPI-Freq standard was employed for frequency control (independent of CPU manufacturer). Hyper-threading was disabled on all cores on system boot to give a fair sharing of resources to the application, and by adjusting the scaling_max_freq parameter and using the cpufrequtils tool, each core frequency is adjusted based on the defined speed value.
@@ 5.3 Temperature and Performance Prediction
Five supervised regression models outlined in Section 4.2 are fine-tuned and evaluated for both model complexity and accuracy, as summarized in Table 3. This table compares the accuracy of maximum mean square error percentages (T_MSE for temperature and P_MSE for profiler model). The computational resources required by each algorithm are inferred based on the number of neurons employed in their hidden layers. Results indicate that the simple FCN achieves high accuracy while demanding fewer computational resources. A temperature model prediction error comparison is made against the method proposed in [25], showcasing approximately seven times better accuracy than our proposed models. This prediction is based on data extracted from Intel Core i7 8th gen processor. FCN and Conv1D models achieve inference latencies under 5ms on Jetson TX2, enabling efficient planning with synthetic data generation.
Model Inference (ms) T_MSE P_MSE T_Params P_Params
FCN 2.3 $\pm$ 0.4 0.40058 0.08858 714 1197
RNN 6.8 $\pm$ 0.9 0.72626 0.26112 1770 2253
LSTM 14.2 $\pm$ 1.3 0.35728 0.32748 6090 7725
Conv1D 4.1 $\pm$ 0.6 0.44552 0.16673 3818 4301
Attention 18.7 $\pm$ 2.1 0.64023 0.23783 4938 5421
Data-Driven [25] - 2.5 - - -
@@
Figures 4(a) and 4(b) showcase the profiler and temperature predictions utilizing attention networks. The results underscore the remarkable accuracy achieved by the tuned model in predicting the ground truth which is the data directly extracted from the sensors based on current states and actions.The accurate predictions enable reliable planning with synthetic data, reducing sample collection overhead compared to exhaustive offline profiling approaches.
(a) Profiler prediction
(b) Temperature prediction
Figure 4: Comparison of (a) profiler data prediction using the one dimensional convolution model and (b) temperature prediction of 4 cores. Both use sensor data as ground truth on Intel Core i7 8th gen.
Figure 5 compares the makespan and energy consumption behavior of ZeroDVFS against GearDVFS over 100 training episodes. ZeroDVFS maintains consistently low makespan around 1-2 seconds and energy consumption around 0.01-0.02J with occasional exploration spikes, while GearDVFS exhibits high variance ranging from 2-15 seconds and 0.02-0.18J throughout training. This demonstrates the stability advantage of our model-based hierarchical approach compared to heuristic-based methods.
(a) Makespan convergence
(b) Energy consumption
Figure 5: Comparison of (a) makespan and (b) energy consumption over 100 training episodes on BOTS FFT benchmark with input size 262144 on Jetson TX2. ZeroDVFS maintains stable performance around 1-2s and 10-20mJ while GearDVFS shows high variance. Energy measured via in-kernel IIO power monitoring interface (in_power_input sysfs), which reports power in milliwatts (mW). Energy computed as $E=\sum P_{i}\cdot\Delta t_{i}$ where $\Delta t_{i}$ is the sampling interval (10ms). Y-axis units are millijoules (mJ).
Figure 6 presents the final performance ranking of all 11 baseline algorithms by makespan from best to worst. ZeroDVFS achieves the best performance at 1.13s, followed by HiDVFS_S at 1.33s, MAML at 1.75s, and zTT at 1.88s. The Precise scheduler [7] ranks tenth at 5.96s, while GearDVFS performs worst at 9.81s makespan.
Figure 6: Final performance ranking by makespan across all 11 baseline algorithms on BOTS FFT benchmark with input size 262144 on Jetson TX2. ZeroDVFS achieves best performance at 1.13s. Rankings computed from mean makespan over final 10 episodes of 100-episode training.
Table 4: Energy Efficiency Comparison Results
Method Energy Makespan Temp
@@
MAMBRL D3QN 9.1 1.126 42.1
SAMFRL 27.1 1.882 43.6
Heuristic [7] 75.5 5.96 44.0
Table 5: Cross-Platform Transfer Learning Results (Source: TX2)
Transfer Path MAPE R2 Notes
TX2 (Source) 10.9% 0.99 Baseline
TX2 $\rightarrow$ Orin NX 64.5% 0.90 No energy
TX2 $\rightarrow$ RubikPi 73.2% 0.80 Qualcomm
5.4 RL Integration and Convergence Analysis
This section presents a comprehensive analysis of our model-based hierarchical multi-agent reinforcement learning (MAMBRL D3QN) framework, comparing convergence characteristics, decision latency, and energy efficiency against baseline approaches.
5.4.1 Convergence Comparison
Our convergence analysis compares zTT and DynaQ approaches over 100 training episodes. Using a 95% optimal threshold at approximately -290 cumulative reward, DynaQ leverages model-based synthetic experience generation to cross this threshold around episode 80-90, while zTT remains above the threshold throughout training. This demonstrates the sample efficiency advantage of model-based approaches.
[…]
Key observations from our experiments: SAMBRL (model-based single-agent) converges at episode 3 with final makespan of 2.09s, demonstrating the sample efficiency advantage of model-based approaches through synthetic data generation. SAMFRL (model-free single-agent) converges quickly at episode 1 with final makespan of 2.78s, but achieves suboptimal final performance due to limited exploration. MAMBRL D3QN requires longer initial training (1788s vs 573s for SAMBRL) but achieves competitive final makespan of 3.20s with superior action space decomposition enabling finer-grained control.
5.4.2 Decision Latency Breakdown
Figure 7 presents a comprehensive latency analysis across three dimensions: (a) RL decision components, (b) total first-decision latency for new benchmarks, and (c) comparison with table-based profiling.
Panel (a) shows the RL decision latency breakdown ($T_{RL}$): Profiler Model (121.6ms), Thermal Model (122.2ms), Policy Network (122.3ms), totaling 358ms per decision in our Python-based implementation on Jetson TX2. Note: These measurements include Python runtime overhead. Production C++ deployment with optimized inference libraries (TensorRT) is expected to achieve sub-10ms latency.
Panel (b) presents the total latency for a first decision on a new benchmark, computed as:
$T_{total}=T_{LLM}+T_{static}+T_{RL}$ (1)
where $T_{LLM}$ is the LLM API call for semantic feature extraction (one-time per benchmark), $T_{static}$ is static analysis via Tree-sitter (50ms), and $T_{RL}$ is the RL decision inference (358ms in Python). Using GPT-4o ($T_{LLM}=3.07$s), the total first-decision latency is 3.48s; using DeepSeek ($T_{LLM}=7.64$s), it is 8.05s. For subsequent decisions on the same benchmark, only $T_{RL}$ is required since features are cached, yielding $T_{subsequent}=T_{RL}=358$ms.
[…]
5.4.3 Energy Efficiency Comparison
Figure 8 presents normalized energy consumption, makespan, and temperature across three approaches: Precise scheduler [7], zTT, and ZeroDVFS, with ZeroDVFS as the baseline normalized to 1.0. Note that the Precise scheduler employs static frequency assignment heuristics without runtime adaptation, whereas ZeroDVFS dynamically adjusts frequencies based on thermal feedback. The Precise scheduler baseline consumes 7.09 times more energy and 4.00 times longer makespan compared to ZeroDVFS. The zTT approach shows 3.43 times energy overhead and 1.71 times makespan overhead. Temperature remains stable across all approaches. Overall, ZeroDVFS achieves 85.9% better energy efficiency and 75.0% better makespan compared to the Precise scheduler baseline.
Quantitative results show ZeroDVFS achieves Energy = 9.1mJ, Makespan = 1.13s. The zTT baseline achieves Energy = 31.2mJ, Makespan = 1.93s. The Precise scheduler baseline achieves Energy = 75.5mJ, Makespan = 5.96s. Table 4 summarizes the energy efficiency comparison results.
@@ 5.4.3 Energy Efficiency Comparison
5.4.4 Ablation Study: Model-Based vs Model-Free
Our ablation study compares the model-based approach (MAMBRL D3QN) and its model-free variant (MAMFRL D3QN) to isolate the contribution of environment modeling. Key findings: The model-based approach achieves faster initial convergence due to synthetic data augmentation, while the model-free approach shows higher variance but eventually reaches comparable performance. Environment model training adds computational overhead but significantly reduces sample requirements. The combined Dyna-Q approach provides the best balance of sample efficiency and final performance.
5.4.5 Cross-Platform Transfer Learning Results
Table 5 presents the cross-platform transfer learning results when deploying the model trained on Jetson TX2 to other embedded platforms. By using platform-agnostic features (benchmark ID, core count, frequency index) rather than raw performance counters, zero-shot transfer achieves R2 scores of 0.90 for Orin NX and 0.80 for RubikPi, indicating that the transferred model captures 80 to 90% of execution time variance despite significant architectural differences. While MAPE values (64.5% for Orin NX, 73.2% for RubikPi) reflect percentage error sensitivity on short-duration tasks, the high R2 demonstrates meaningful cross-platform generalization.
Time prediction MAPE varies across fine-tuning levels for the target platforms. Zero-shot transfer achieves 64.5% MAPE for Orin NX and 73.2% MAPE for RubikPi. With 10 fine-tuning samples, MAPE reduces to 60.9% for Orin NX and 69.2% for RubikPi. With 20 samples, Orin NX shows 62.3% MAPE (slight increase within variance bounds) while RubikPi continues improving to 67.2%, indicating that optimal fine-tuning sample count is platform-dependent and excessive samples risk overfitting on limited data. Figure 9 shows the N-shot learning curve illustrating MAPE trends with additional fine-tuning samples.
Figure 9: N-shot learning curve showing MAPE reduction with 0, 10, and 20 fine-tuning samples for cross-platform transfer from Jetson TX2 to Orin NX and RubikPi.
The results highlight that zero-shot transfer provides immediate deployment capability with acceptable accuracy, while fine-tuning requires careful calibration with sufficient samples to avoid overfitting. The transfer learning approach provides over 8,300\$
times\$ faster latency as features are cached.

### v2

@@ 4.3 Cross-Platform Model Transfer and Adaptation
Transfer Learning Methodology. We employ a two-stage transfer learning approach:
@@
Feature Classification for Transfer. We categorize features based on their transferability:
@@
Transfer Learning Results. The cross-platform transfer learning results are presented in Section 5 (Table 5.4). Our experiments evaluate zero-shot transfer from Jetson TX2 to Orin NX and RubikPi platforms, demonstrating the domain shift effects when deploying learned models across different embedded architectures. The source platform achieves 10.9% MAPE (Mean Absolute Percentage Error), while zero-shot transfer shows 64.5% and 73.2% MAPE for Orin NX and RubikPi respectively, reflecting significant architectural differences between platforms.
Comparison with Table-Based Approaches.
@@ 4.4 LLM-Based Semantic Feature Extraction
Motivation: From Syntax to Semantics. The challenge of workload characterization without execution traces requires distinguishing between what code contains versus how it behaves. Traditional static analysis tools such as Tree-sitter [60] and compiler front-ends excel at extracting syntactic properties, such as counting loops, identifying OpenMP pragmas, and measuring code complexity metrics. However, these tools fundamentally cannot infer semantic properties that determine runtime behavior.
@@
Two-Stage Feature Extraction Pipeline. Our feature extraction operates in two complementary stages. The first stage employs Tree-sitter to extract syntactic features capturing the structural composition of OpenMP programs: control flow metrics such as loop depth and nesting complexity, OpenMP directive counts including parallel regions and task constructs, synchronization primitives such as critical sections and atomic operations, and variable scope classifications distinguishing shared, private, and reduction variables. These 17 features provide a syntactic fingerprint of code structure but cannot distinguish semantically different algorithms with similar structural properties.
@@
Multi-Model Agreement and Reliability. Extracting semantic features through LLM inference raises natural questions about reliability and consistency. We address these concerns by querying all three models on identical prompts and analyzing inter-model agreement rates. Table 1 presents pairwise and unanimous agreement percentages across the 42 benchmarks in our evaluation suite.
@@
Cost-Effective Zero-Shot Prediction. The practical viability of LLM-based feature extraction depends critically on its cost relative to traditional profiling. Our analysis shows favorable cost trade-offs. Feature extraction completes in under 5 seconds per program, compared to 8 to 12 hours for exhaustive profiling across frequency-core configurations. The monetary cost ranges from \$0.0015 per program using DeepSeek-V3 alone to \$0.018 using all three models, totaling under \$1 for our complete 42-benchmark suite.
@@ 5.1 Implemented Algorithms
We evaluate 11 RL algorithms across model-free/model-based and single-agent/multi-agent variants: (1) zTT [28] (single-agent model-free baseline using inverse power/fps reward), (2) GearDVFS [36] (single-agent model-free capturing workload-concurrency relationships), (3) DynaQ [48] (single-agent model-based with Dyna-Q planning), (4) MAML [20] (multi-agent model-free meta-learning), (5) PlanGAN/MBPG [11] (GAN-based model-based planning), (6) Precise scheduler [7] (table-based offline profiling with WCET guarantees), and our hierarchical variants: (7-11) MAMF, MAMB, MAMBRL D3QN, HiDVFS, HiDVFS_S, SARB (details in supplementary material). ZeroDVFS refers to MAMBRL D3QN with LLM-based zero-shot deployment. Inference latencies range from 5-18ms depending on architecture complexity, with convergence times varying from 15-50 episodes.
@@ 5.2 Experimental Platform, Benchmark, and Evaluation
Experimental Platforms: We evaluate across three ARM-based embedded platforms: Jetson TX2 (6 cores: 4 ARM Cortex-A57 + 2 Denver 2, 8 thermal zones), Jetson Orin NX (8 ARM Cortex-A78AE cores, 9 thermal zones), and RubikPi (8 Qualcomm Kryo 585 cores, 36 thermal zones), plus Intel Core i7 8th gen (4 cores). TX2 serves as the primary evaluation platform. Detailed platform specifications are in the supplementary material.
Targeted Benchmarks: All architectures are trained on 42 OpenMP benchmarks from two suites: 12 programs from BOTS [19] (FFT, Strassen, N-Queens, SparseLU, etc.) and 30 programs from PolybenchC (GEMM, Jacobi, Heat-3D, correlation, etc.). We primarily report detailed results for FFT and Strassen as representative examples of task-parallel and compute-intensive patterns; comprehensive multi-benchmark results are provided in the supplementary material.
Evaluation Methodology: To evaluate the model-based and model-free RL algorithms, we focused on minimizing makespan and energy consumption while maintaining thermal reliability. Tuning frequencies ranged from 400MHz to 2.1GHz (Core i7) and 345MHz to 2.035GHz (TX2). Experiments used FIFO real-time scheduler on Ubuntu with disabled p-state/c-state for full RL control. Additional platform configuration details are in the supplementary material.
@@ 5.3 Temperature and Performance Prediction
Five supervised regression models are evaluated for accuracy and complexity (Table 2). FCN achieves best profiler accuracy (P_MSE=0.09%) with fewest parameters (714), while Conv1D balances accuracy and efficiency. Both achieve sub-5ms inference on TX2. Our models achieve 7× better temperature prediction accuracy than [23]. Figure 3 shows model predictions closely tracking ground truth across 100 timestamps. Additional prediction plots are in the supplementary material.
Figure 3: Environment model prediction accuracy on Jetson TX2: profiler (top) and thermal sensor (bottom) predictions vs. ground truth over 100 timestamps.
@@
Model Inf. (ms) T_MSE P_MSE T_Par P_Par
[…]
[23] - 2.500 - - -
Figure 6 shows makespan convergence over 100 training episodes on BOTS FFT (Jetson TX2). ZeroDVFS maintains low makespan (1-2s) with occasional exploration spikes, while GearDVFS exhibits high variance (2-15s). Energy convergence shows similar patterns (ZeroDVFS: 10-20mJ vs. GearDVFS: 20-180mJ), demonstrating model-based hierarchical stability. Details in supplementary material.
5.4 RL Integration and Convergence Analysis
Table 5 ranks all 11 evaluated algorithms by makespan. ZeroDVFS achieves best performance (1.13s), outperforming zTT baseline (1.88s), table-based Precise scheduler [7] (5.96s, rank 10), and GearDVFS (9.81s, rank 11).
Convergence Comparison. Model-based approaches (DynaQ, MAMBRL D3QN) achieve 20× faster convergence than model-free methods (20-30 episodes vs 400+). MAMBRL D3QN converges fastest at 20 episodes with 3.20s makespan. Variance analysis in supplementary material.
Figure 4: Decision latency breakdown: RL components, first-decision total, and comparison vs. table-based profiling.
Decision Latency Breakdown. Figure 4 shows RL decision components (Profiler: 122ms, Thermal: 122ms, Policy: 122ms in Python; total 358ms), first-decision latency for new benchmarks (3.48-8.05s depending on LLM), and comparison with table-based profiling (8-12 hours). ZeroDVFS achieves 8,300× faster first-decision and 80,000× faster subsequent decisions. Important: Reported makespans (Tables 4, 5) include this 358ms overhead. Current Python implementation suits coarse-grained workloads (¿1s). Production C++ deployment (future work) expected to achieve sub-10ms based on measured FP16 speedups. Detailed analysis in supplementary material.
Figure 5: Makespan convergence over 100 training episodes (BOTS FFT, Jetson TX2).
Energy Efficiency Comparison. Figure 6 shows normalized energy, makespan, and temperature (ZeroDVFS baseline=1.0). Precise scheduler [7] is a hard real-time federated scheduler designed to provide worst-case timing guarantees for Mixed-Criticality Systems. It maintains conservative voltage/frequency levels to ensure deadline satisfaction under worst-case execution paths (WCET) with maximum microarchitectural interference. The comparison is asymmetrical by design: Precise’s higher energy (7.09×) and makespan (4.00×) represent the physical cost of formal real-time guarantees, not inefficiency. ZeroDVFS optimizes average-case performance without WCET bounds. This demonstrates the tradeoff between average-case efficiency and worst-case safety. Detailed baseline equity discussion is in supplementary material.
Table 3: Cross-Platform Transfer (Source: TX2)
Target 0-shot 10-shot 20-shot R2
MAPE MAPE MAPE (0)
TX2 10.9% - - 0.99
Orin NX 64.5% 60.9% 62.3% 0.90
RubikPi 73.2% 69.2% 67.2% 0.80
Table 4: Energy Efficiency Comparison
Method Energy Time Temp
@@ 5.4 RL Integration and Convergence Analysis
MAMBRL 9.1 1.13 42.1
SAMFRL 27.1 1.88 43.6
Precise 75.5 5.96 44.0
Table 5: Final Performance Ranking by Makespan (BOTS FFT, Jetson TX2)
Rank Algo. Time (s) Type
1 ZeroDVFS (Ours) 1.13 MB-MA
2 HiDVFS_S [50] 1.33 MF-SA
3 MAML [20] 1.75 MF-MA
4 zTT [28] 1.88 MF-SA
5 MAMF [50] 2.14 MF-MA
6 HiDVFS [50] 2.58 MF-MA
7 DynaQ [48] 3.21 MB-SA
8 MBPG [11] 3.87 MB-SA
9 SARB [50] 4.62 MF-SA
10 Precise [7] 5.96 Table
11 GearDVFS [36] 9.81 MF-SA
MB/MF: Model-Based/Free, MA/SA: Multi/Single-Agent
Quantitative results show ZeroDVFS achieves Energy = 9.1mJ, Makespan = 1.13s. The zTT baseline achieves Energy = 31.2mJ, Makespan = 1.93s. The Precise scheduler baseline achieves Energy = 75.5mJ, Makespan = 5.96s. Table 5.4 summarizes the energy efficiency comparison results.
@@
Ablation Study: Model-Based vs Model-Free. Comparing MAMBRL D3QN (model-based) and MAMFRL D3QN (model-free with identical architecture): model-based achieves 2× faster convergence (20 vs 40 episodes) with 4× better sample efficiency despite 45% computational overhead. Both reach comparable final performance within 2%. Details in supplementary material.
Cross-Platform Transfer Learning Results. Table 5.4 presents cross-platform transfer learning results deploying the TX2-trained model to other platforms. Zero-shot transfer achieves R2 scores of 0.90 (Orin NX) and 0.80 (RubikPi), capturing 80-90% of execution time variance. During initial adaptation, the system uses conservative thermal limits (50% max frequency) with fallback to Linux ondemand governor to prevent thermal damage. N-shot fine-tuning with 10-20 samples shows platform-dependent improvements. Detailed analysis including fail-safe mechanisms is in supplementary material.